The sum of ages of mother & father & son is 96 years after 8 years mother will be three times as old as her son & father will be as old as mother & son together find the present age of mother
Answer by ankor@dixie-net.com(22740)

The sum of ages of mother & father & son is 96 years
m + f + s = 96
after 8 years mother will be three times as old as her son
m + 8 = 3(s+8)
m + 8 = 3s + 24
m - 3s = 24 - 8
m - 3s = 16
:
& father will be as old as mother & son together
f + 8 = (m+8 + s+8)
f = m + s + 16 - 8
f = m + s + 8
f - m - s = 8
Add the above equation to the first equation
f - m - s = 8
f + m + s = 96
---------------Addition eliminates m and s, find f
2f = 104
f = 104/2
f = 52 yrs is Father's age
:
Back to the first equation, replace f with 52
52 + m + s = 96
m + s = 96 - 52
m + s = 44
Subtract the 2nd equation from the above equation
m + s = 44
m - 3s =16
-----------------
4s = 60
s = 60/4
s = 15 yrs is son's age
:
find the present age of mother
The first equation, replace f and s
52 + m + 15 = 96
m = 96 - 67
m = 29 yrs is Mom's age
:
;
See if that checks out in the statement
"After 8 years...father will be as old as mother & son together "
52 + 8 = 29 + 8 + 15 + 8
60 = 37 + 23